please find below a short summary of the activities of the the related to
the AO system.
People at LBTO:
A. Riccardi and M.Xompero (VNC+Skype)
The shell was successfully set at the first attempt, at 02:02UT. Due to
unfavourable weather conditions (sky a bit cloudy and strong wind), the dome
will not be opened all the night long. Hence, we performed functionality
tests on the AdSec
As a first test, we moved the telescope to different elevations and set the
shell, recording the result of the load_shape procedure (i.e. the actuators
forces and positions after the application of the given command). As a
second test, we set the shell while the telescope was pointing the zenith
and moved it to 30° elevation, saving actuators readings. Such test was
performed twice, both at fast and slow telescope speed).
During the test, the shell has been regularly reset in order to apply the
wanted shape; while in rest position, the gap inspection revealed that the
"ear" on the shell side disappeared (i.e. the gap is at maximum 39um).
A software error was found in the routine that automatically correct for the
astigmatism, while the telescope changes elevation: at 90° a huge
astigmatism was applied, then removed, as a consequence of a wrong
elevationreading. The bug has been fixed.
With the exception of a sigle event (the AdSecArbitrator
no problem was registered during the night.
A more exhaustive log may be found here: TelLogLBT672a20100401
A spreadsheet with system temperatures and additional test informations
(i.e. telemetry timestamps) has been saved. (see log_temp.xls
- 02 Apr 2010